# Graph-Theoretic Analysis of $n$-Replica Time Evolution in the Brownian Gaussian Unitary Ensemble

### v1

@@ 2.1 Notations of a Graph
Here, $c[\sigma]$ is the permutation on the contour labels on the right-hand side, $c_{i\bar{j}}$ represents the pair between the left sites $i,\bar{j}$ and the right sites $\sigma(i)^{\prime},\sigma(\bar{j})^{\prime}$. To make the complicated notations clear, we can take a simple example $c_{1\bar{1}}c[X_{12}\otimes X_{\bar{1}\bar{2}}]=\delta_{I_{1}J_{1}}\delta_{I_{2}^{\prime}J_{2}^{\prime}}\delta_{I_{2}I_{1}^{\prime}}\delta_{J_{2}J_{1}^{\prime}}$ and show how the graphs work as follows:
@@
The filled dots are the sites on the contour $H$ and the hollow circles are those on the dual contour $H^{*}$. And we can directly find how the swapping and connecting operations work in the first two steps. The graph is finally denoted as the R.H.S. of (17).
The notations presented in (16) for the graphs make it much easier to calculate the contraction (product) of two graphs and study the action of $\mathcal{L}_{n}$ on a certain graph, as we will see later in this section.
Symmetries of a Graph
Before examining the action of $P_{i\bar{j}}$ and $X_{ij},X_{\bar{i}\bar{j}}$ on a graph $\mathsf{F}$, we note that the notations presented above have some redundancy. For a graph $\mathsf{F}=\prod_{i=1}^{p}c_{a_{i}\bar{b}_{i}}c[\sigma]$, the vertices for each connected pair $(a_{i},\bar{b}_{i})$ are identical. To simplify it, we define the action of a permutation on the graph as
@@
If a permutation $\sigma^{\prime}=\sigma^{\prime}_{H}\otimes\sigma^{\prime}_{H^{*}}$ is an element of the group $G_{n}(\boldsymbol{a},\bar{\boldsymbol{b}})$ generated by $\{X_{a_{i}a_{j}}\otimes X_{\bar{b}_{i}\bar{b}_{j}}\}_{i,j=1,2,\ldots,p}$, then its action is trivial: $\sigma^{\prime}*\mathsf{F}=\prod_{i=1}^{p}c_{a_{i}\bar{b_{i}}}c[\sigma\sigma^{\prime}]=\mathsf{F}$. It can be observed that $G_{n}(\boldsymbol{a},\bar{\boldsymbol{b}})$ resembles the gauge group in quantum field theory and is isomorphic to the permutation group $S_{p}$.
@@
$N_{n}$ $=\sum_{\boldsymbol{a},\bar{\boldsymbol{b}}}{|S_{n}\otimes S_{n}|\over|G_{n}(\boldsymbol{a},\bar{\boldsymbol{b}})|}=\sum_{p=0}^{n}\binom{n}{p}^{2}p!\frac{\left(n!\right)^{2}}{p!}=\sum_{p=0}^{n}\left(n!\binom{n}{p}\right)^{2}\equiv\sum_{p=0}^{n}N_{n,p}\ .$ (19)
One can utilize the individual graphs as a basis, then $\mathcal{L}_{n}\mathsf{F}_{i}=\sum_{j=1}^{N_{n}}\widetilde{M}_{ji}\mathsf{F}_{j}$. However, as indicated in (19), the dimension of $\widetilde{M}$ grows too quickly with $n$. A clever approach is to group these individual graphs into categories $\mathsf{F}_{\alpha}=\sum_{i=1}^{N_{n;\alpha}}\mathsf{F}_{\alpha,i}$ such that
$\mathcal{L}_{n}\mathsf{F}_{\alpha}=\sum_{\beta=1}^{N_{n}^{\text{ca}}}M_{\beta\alpha}\mathsf{F}_{\beta}\ .$ (20)
In this paper, we further require that $\mathsf{F}_{\alpha}\cap\mathsf{F}_{\beta}=\emptyset$ for all $\alpha\neq\beta$. The dimension of the transfer matrix $M$, denoted as $N_{n}^{\text{ca}}$, can be significantly smaller than $N_{n}$, as illustrated in Table 1.
2.2 Operating $\mathcal{L}_{n}$ on an Individual Graph
Figure 1: Four non-trivial cases of $P_{i\bar{j}}\mathsf{F}$.
Since $\mathcal{L}_{n}$ consists of $P_{i\bar{j}}$ and $X_{ij},X_{\bar{i}\bar{j}}$, we first evaluate their action on an individual graph $\mathsf{F}=\prod_{i=1}^{p}c_{a_{i}\bar{b}_{i}}c[\sigma]$. We denote the set of paired sites as $C(\mathsf{F})=\{(a_{i},\bar{b}_{i})\}_{i=1}^{p}$. For later reference, we also define the set of all possible paired sites between different contours as $\mathcal{C}$ and those on the same contour as $\mathcal{E}$ and $\bar{\mathcal{E}}$, respectively:
@@ 2.2 Operating $\mathcal{L}_{n}$ on an Individual Graph
Operation of $\mathcal{L}_{n}$
Next, we can study the operation of $\mathcal{L}_{n}$. Notice that we have the condition $\bar{u}\ \cap\ C[\mathsf{F}]=\emptyset$. Hence the permutations $X_{u}$ and $X_{\bar{u}}$ act trivially on $C(\mathsf{F})$. Thus, we ultimately find a simple relation
@@
where $|C[\mathsf{F}]|$ counts the number of pairs, and $\mathsf{F}[\sigma\sigma^{\prime}]\equiv\prod_{i=1}^{p}c_{a_{i}\bar{b}_{i}}c[\sigma\sigma^{\prime}]$. The definitions of $u$ and $\bar{u}$ are encoded in the summation expressions. For example, in the term $\sum_{u\cap C[\mathsf{F}]=\emptyset}c_{u}\mathsf{F}$, $u$ can be considered as a pair in $\mathcal{C}$ that is excluded by $C(\mathsf{F})$. Meanwhile, in the term $\sum_{u\cap C[\mathsf{F}]=\emptyset}\mathsf{F}[\sigma X_{u}]$, $u$ can be regarded as an element of $\mathcal{E}$ that is excluded by $C(\mathsf{F})$. Here we choose the set of $u$ to meet those conditions of $u\cap C[\mathsf{F}]=\emptyset$ from all the summations, which requires flattening $C(\mathsf{F})$ into a list of site labels and find those $u$ that are available. For example, taking $u=\{(1,\bar{1})\},C(\mathsf{F})=\{(1,\bar{2}),(2,\bar{3})\}$, then
$u\cap C(\mathsf{F})=\{1,\bar{1}\}\cap\{1,\bar{2},2,\bar{3}\}=\{1\}\not=\emptyset\ ,$ (24)
which shows that $u=\{(1,\bar{1})\}$ is not available.
@@
$\mathsf{F}_{0}\equiv c[I],\mathsf{F}_{1}\equiv\sum_{u}c_{u}c[I],\mathsf{F}_{0,X+\bar{X}}\equiv\sum_{e}c[X_{e}]+\sum_{\bar{e}}c[X_{\bar{e}}]\ .$ (25)
Then we apply $\mathcal{L}_{n}$ to the three initial categories and their descendants until a complete basis is obtained. After that, we obtain the expression for the transfer matrix $M$ in (20), which allows us to solve $\mathcal{U}_{n}(t)$ using (5). We will illustrate the basic idea with examples for $n=2$ and $n=3$ in the next two sections.
Figure 2: Two non-trivial cases of $X_{ij}\mathsf{F}$.
@@ 3 Time Evolution for $n=2$
As a warm-up, we consider $n=2$, which gives us four sites: $\{1,\bar{1},2,\bar{2}\}$. The permutation group is $S_{2}\otimes S_{2}$, where $S_{2}=\{I,X_{12}\}$. Then
$\mathcal{C}$ =$=\left\{\left(1,\bar{1}\right),\left(1,\bar{2}\right),\left(2,\bar{1}\right),\left(2,\bar{2}\right)\right\}\ ,$ (26)
@@
The number of individual graphs is $24$, while the number of operators ($P_{i\bar{j}},X_{ij},X_{\bar{i}\bar{j}}$) is $6$. Therefore, in principle, one needs to calculate $6\times 24=168$ graph contractions when calculating $M$ as defined earlier. Fortunately, (23) enables us to perform a simple algebraic calculation without drawing any graphs.
@@ 3.1 Counting Graph Categories
As we have discussed earlier, we will find the appropriate graph groups by applying $\mathcal{L}_{n}$ to the three initial categories $\mathsf{F}_{0},\mathsf{F}_{1},\mathsf{F}_{0,X+\bar{X}}$ and their descendants until no new graph category is generated. In the case of $n=2$, we will perform $\mathcal{L}_{2}$ on the initial sets as follows.
@@
  • $\mathcal{L}_{2}\mathsf{F}_{0}$: The operation of $\mathcal{L}_{2}$ on $\mathsf{F}_{0}$ is trivial, since $\mathsf{F}_{0}$ is a trivial set
@@
    $\mathcal{L}_{2}\mathsf{F}_{1}=\left(w+J\right)\mathsf{F}_{1}+\frac{2J}{D}\mathsf{F}_{2}\ ,$ (30)
@@
    then we can define $\mathsf{F}_{1,X+\bar{X}}\equiv\sum_{u}c_{u}\mathsf{F}_{0,X+\bar{X}}$. Notice that
@@
We find that three new graph categories, $\mathsf{F}_{2},\mathsf{F}_{1,X+\bar{X}},\mathsf{F}_{0,X\bar{X}}$, emerge after applying $\mathcal{L}_{2}$ on initial sets, which is the first order calculations. Next, we consider the second order, which means the operation of $\mathcal{L}_{2}$ on these new categories. For clarity, we restate the definitions here
$\mathsf{F}_{2}=c_{1\bar{1}}c_{2\bar{2}}c[I]+c_{1\bar{2}}c_{2\bar{1}}c[I]\ ,$ (35)
$\mathsf{F}_{1,X+\bar{X}}=\sum_{u}c_{u}\mathsf{F}_{0,X+\bar{X}}\ ,$ (36)
$\mathsf{F}_{0,X\bar{X}}=c[X_{12}X_{\bar{1}\bar{2}}]\ .$ (37)
@@
  • $\mathcal{L}_{2}\mathsf{F}_{2}$: Notice that only the first term in (23) survives. The operation of $\mathcal{L}_{2}$ on $\mathsf{F}_{2}$ is straightforward to obtain
@@
  • $\mathcal{L}_{2}\mathsf{F}_{1,X+\bar{X}}$: Using (23), we have
@@
  • $\mathcal{L}_{2}\mathsf{F}_{0,X\bar{X}}$: Using (23), we have
@@
Here, we find two new graph categories, $\mathsf{F}_{2,X+\bar{X}}$ and $\mathsf{F}_{1,X\bar{X}}$. We will then study the third order calculations by operating $\mathcal{L}_{2}$ on these categories
$\mathsf{F}_{2,X+\bar{X}}=c_{1\bar{1}}c_{2\bar{2}}c[X_{12}]+c_{1\bar{2}}c_{2\bar{1}}c[X_{12}]\ ,$ (43)
$\mathsf{F}_{1,X\bar{X}}=\sum_{u}c_{u}c[X_{12}X_{\bar{1}\bar{2}}]\ .$ (44)
@@
The calculations are straightforward by using (23)
  • $\mathcal{L}_{2}\mathsf{F}_{2,X+\bar{X}}$:
@@
  • $\mathcal{L}_{2}\mathsf{F}_{1,X\bar{X}}$:
@@
    $=\left(w+J\right)\mathsf{F}_{1,X\bar{X}}+\frac{2J}{D}\mathsf{F}_{2}\ ,$
@@
Rank $r$ New Categories #$\#\mathsf{F}$
@@ 4 $\emptyset$ 0
So far, we find that no new categories emerge. In summary, we list the total of 8 categories in Table 2.
@@ 3.2 Observables
$\begin{bmatrix}f_{1}(t)\\ f_{2}(t)\\ f_{3}(t)\\ f_{4}(t)\\ f_{5}(t)\\ f_{6}(t)\\ f_{7}(t)\\ f_{8}(t)\end{bmatrix}=\begin{bmatrix}0&0&\frac{1}{4}&\frac{1}{2}&\frac{1}{4}\\ 0&\frac{D^{2}-2}{D\left(D^{2}-4\right)}&\frac{-1}{4(D-2)}&\frac{-1}{2D}&\frac{-1}{4(D+2)}\\ 0&0&-\frac{1}{4}&0&\frac{1}{4}\\ 0&\frac{-1}{D^{2}-4}&\frac{1}{4(D-2)}&0&\frac{-1}{4(D+2)}\\ \frac{1}{D^{2}-1}&\frac{-2}{D^{2}-4}&\frac{1}{2(D-1)(D-2)}&0&\frac{1}{2(D+1)(D+2)}\\ 0&0&\frac{1}{4}&-\frac{1}{2}&\frac{1}{4}\\ \frac{-1}{D^{3}-D}&\frac{4}{D(D^{2}-4)}&\frac{-1}{2(D-1)(D-2)}&0&\frac{1}{2(D+1)(D+2)}\\ 0&\frac{2}{D(D^{2}-4)}&\frac{-1}{4(D-2)}&\frac{1}{2D}&\frac{-1}{4(D+2)}\end{bmatrix}\times\begin{bmatrix}1\\ e^{-Jt}\\ e^{-(2-2D)^{-1}Jt}\\ e^{-2Jt}\\ e^{-(2+2D)^{-1}Jt}\end{bmatrix}\ .$ (58)
Generally, we can calculate OTOC by using the expression of (57)
$\text{Tr}\left(O_{1}(t)O_{2}O_{3}(t)O_{4}\right)=\leavevmode\hbox{$\vbox{\hbox{\resizebox{}{}{{\leavevmode\hbox{\set@color{}}}}}}$}\equiv\sum_{a=1}^{8}f_{a}(t)\times\left(e^{-\text{i}\mathsf{E}t}\mathsf{F}_{a}\right)_{O_{1},O_{2},O_{3},O_{4}}\ .$ (59)
A special case would be the noise-free case, which means that the noise strength $J$ vanishes. We expect that the time evolution will be simplified into a unitary one. Note that from (58) we can directly find that all $f_{a}(t)$ vanishes except $f_{1}(t)=1$. Hence
@@
which is just the unitary time evolution for a closed system and finishes our consistency check. Thus, we will be more confident to use (59) to calculate the fluctuations of two-point functions, spectral form factor (SFF) and other observables.
To end this section, we would mention that there is an alternative method Guo:2024zmr to evaluate $\mathcal{U}_{2}(t)$ by observing that
@@
We can always decompose a wavefunction $\psi=\psi_{i_{1}i_{2}}^{j_{1}j_{2}}\ket{i_{1},j_{1},i_{2},j_{2}}_{1\bar{1}2\bar{2}}$ into two singlets $\mathbf{1}$, four adjoints $\mathbf{D}^{2}-\mathbf{1}$, an antisymmetric traceless $\mathbf{A}$, two mixed $\mathbf{M}$ and a symmetric traceless $\mathbf{S}$ rank $(2,2)$ irreducible representations of $\text{SU}(D)$, with all the coefficients. Thus, we can also do the exponentiation on this basis. However, notice that the dimension of the representation is $10$, which is larger than the number of independent graph categories $8$, indicating that there are still some redundancies in the Young tableaux method. This might cause a longer processing time in numerical calculations.

### v2

@@ 2.1 Notation for a graph
Here, $c[\sigma]$ means to act the permutation on the contour labels on the right-hand side, $c_{i\bar{j}}$ means to connect the left points $i,\bar{j}$ and right points $\sigma(i)^{\prime},\sigma(\bar{j})^{\prime}$; finally, we connect the remaining points $i,\sigma(i)^{\prime}$ and $j,\sigma(\bar{j})^{\prime}$. Taking $c_{1\bar{1}}c[X_{12}\otimes X_{\bar{1}\bar{2}}]=\delta_{I_{1}J_{1}}\delta_{I_{2}^{\prime}J_{2}^{\prime}}\delta_{I_{2}I_{1}^{\prime}}\delta_{J_{2}J_{1}^{\prime}}$ as an example:
@@
Since we use $I$ to indicate the indices of the contours $1,2$ and $J$ for $\bar{1},\bar{2}$, we do not distinguish $J_{1}$ and $J_{\bar{1}}$. In this paper, we use black disks to label contour vertices and circles to label the dual vertices. For simplicity, we will not draw the channel labels $i,\bar{i}$ or the indices $I_{i},J_{i}$ in the graphs.
Moreover, the denotation given in Eq.(15) for the graphs makes it much easier to calculate the contraction (product) of two graphs and study the action of $\mathcal{L}_{n}$ on a certain graph.
Symmetry of a graph
Before examining the action of $P_{i\bar{j}}$ and $X_{ij},X_{\bar{i}\bar{j}}$ on a graph $\mathsf{F}$, we note that the notation presented here contains some redundancy. For a graph $\mathsf{F}=\prod_{i=1}^{p}c_{a_{i}\bar{b}_{i}}c[\sigma]$, the vertices for each connected pair $(a_{i},\bar{b}_{i})$ are identical. To simplify matters, we define the action of a permutation on the graph
@@
If a permutation $\sigma^{\prime}=\sigma^{\prime}_{1}\otimes\sigma^{\prime}_{2}$ is an element of the group $G_{n}(\boldsymbol{a},\bar{\boldsymbol{b}})$ generated by $\{X_{a_{i}a_{j}}\otimes X_{\bar{b}_{i}\bar{b}_{j}}\}_{i,j=1,2,\ldots,p}$, then its action is trivial: $\sigma^{\prime}*\mathsf{F}=\prod_{i=1}^{p}c_{a_{i}\bar{b_{i}}}c[\sigma\sigma^{\prime}]=\mathsf{F}$. It can be observed that $G_{n}(\boldsymbol{a},\bar{\boldsymbol{b}})$ resembles the gauge group in quantum field theory and is isomorphic to the permutation group $S_{p}$.
@@
$N_{n}$ $=\sum_{\boldsymbol{a},\bar{\boldsymbol{b}}}{|S_{n}\otimes S_{n}|\over|G_{n}(\boldsymbol{a},\bar{\boldsymbol{b}})|}=\sum_{p=0}^{n}\binom{n}{p}^{2}p!\frac{\left(n!\right)^{2}}{p!}=\sum_{p=0}^{n}\left(n!\binom{n}{p}\right)^{2}\ .$ (18)
One can utilize the individual graphs as a basis, then $\mathcal{L}_{n}\mathsf{F}_{i}=\sum_{j=1}^{N_{n}}\widetilde{M}_{ji}\mathsf{F}_{j}$. However, as indicated in Eq. (18), the dimension of $\widetilde{M}$ grows too quickly with $n$. A clever approach is to group these individual graphs into categories $\mathsf{F}_{\alpha}=\sum_{i=1}^{N_{n;\alpha}}\mathsf{F}_{\alpha,i}$ such that
$\mathcal{L}_{n}\mathsf{F}_{\alpha}=\sum_{\beta=1}^{N_{n}^{\text{ca}}}M_{\beta\alpha}F_{\beta}\ .$ (19)
In this paper, we further require that $\mathsf{F}_{\alpha}\cap\mathsf{F}_{\beta}=\emptyset$ for all $\alpha\neq\beta$. The dimension of $M_{\beta\alpha}$, denoted as $N_{n}^{\text{ca}}$, can be significantly smaller than $N_{n}$, as illustrated in Table 1.
2.2 Action of $\mathcal{L}_{n}$ on an individual graph
Figure 1: The four non-trivial cases of $P_{i\bar{j}}\mathsf{F}$.
Since $\mathcal{L}_{n}$ consists of $P_{i\bar{j}}$ and $X_{ij},X_{\bar{i}\bar{j}}$, we first evaluate their action on an individual graph $\mathsf{F}=\prod_{i=1}^{p}c_{a_{i}\bar{b}_{i}}c[\sigma]$. We denote the set of pair contours as $C(\mathsf{F})=\{(a_{i},\bar{b}_{i})\}_{i=1}^{p}$. For later reference, we define the set of all possible pair contours as $\mathcal{C}$ and the exchange contours as $\mathcal{E}$ and $\bar{\mathcal{E}}$:
@@ 2.2 Action of $\mathcal{L}_{n}$ on an individual graph
Action of $\mathcal{L}_{n}$
Next, we can study the action of $\mathcal{L}_{n}$. Noting the condition $\bar{u}\cap C[\mathsf{F}]=\emptyset$, the permutations $X_{u}$ and $X_{\bar{u}}$ act trivially on $C(\mathsf{F})$. Thus, we ultimately find a simple relation
@@
where $|C[\mathsf{F}]|$ counts the number of pairs, and $\mathsf{F}[\sigma\sigma^{\prime}]\equiv\prod_{i=1}^{p}c_{a_{i}\bar{b}_{i}}c[\sigma\sigma^{\prime}]$. The definitions of $u$ and $e$ are encoded in the summation expressions. For example, in the term $\sum_{u\cap C[\mathsf{F}]=\emptyset}c_{u}\mathsf{F}$, $u$ can be considered as a pair in $\mathcal{C}$ that is excluded by $C(\mathsf{F})$. Meanwhile, in the term $\sum_{u\cap C[\mathsf{F}]=\emptyset}\mathsf{F}[\sigma X_{u}]$, $u$ can be regarded as an element of $\mathcal{E}$ that is excluded by $C(\mathsf{F})$. There is a commitment to taking the common set $u\cap C[\mathsf{F}]$, which requires flattening $C(\mathsf{F})$ into a list of contour labels. For example, taking $u=\{(1,\bar{1})\},C(\mathsf{F})=\{(1,\bar{2}),(2,\bar{3})\}$, then
$u\cap C(\mathsf{F})=\{1,\bar{1}\}\cap\{1,\bar{2},2,\bar{3}\}=\{1\}\not$$\emptyset\$. (23)
@@
$\mathsf{F}_{0}\equiv c[I],\mathsf{F}_{1}\equiv\sum_{u}c_{u}c[I],\mathsf{F}_{0,X+\bar{X}}\equiv\sum_{e}c[X_{e}]+\sum_{\bar{e}}c[X_{\bar{e}}]\ ,$ (24)
then we apply $\mathcal{L}_{n}$ to the three initial categories and their descendants until a complete basis is obtained. After that, we obtain the expression for $M$ in Eq. (19), which allows us to solve $\mathcal{U}_{n}(t)$ using Eq. (4). We illustrate the basic idea with examples for $n=2$ and $n=3$.
Figure 2: The two non-trivial cases of $X_{ij}\mathsf{F}$.
@@ 3 Time evolution for $n=2$
As a warm-up, we consider $n=2$, which gives us four contours: $\{1,\bar{1},2,\bar{2}\}$. The permutation group is $S_{2}\otimes S_{2}$, where $S_{2}=\{I,X_{12}\}$. Then
$\mathcal{C}$ $=\left\{\left(1,\bar{1}\right),\left(1,\bar{2}\right),\left(2,\bar{1}\right),\left(2,\bar{2}\right)\right\}\ ,$ (25)
@@
The number of individual graphs is $24$, while the number of operators ($P_{i\bar{j}},X_{ij},X_{\bar{i}\bar{j}}$) is $6$. Therefore, in principle, one needs to calculate $6\times 24=168$ graph contractions when evaluating $M$ as defined earlier. Fortunately, Eq. (22) enables us to perform a simple algebraic calculation without the need to draw or recognize any graphs.
@@ 3.1 Counting graph categories
As discussed earlier, we aim to find the appropriate graph groups by applying $\mathcal{L}_{n}$ to the three initial sets $\mathsf{F}_{0},\mathsf{F}_{1},\mathsf{F}_{0,X+\bar{X}}$ and their descendants until no new graph sets are generated.
@@
  • $\mathcal{L}_{2}\mathsf{F}_{0}$: The action of $\mathcal{L}_{2}$ on $\mathsf{F}_{0}$ is trivial
@@
    $\mathcal{L}_{2}\mathsf{F}_{1}=\left(w+J\right)\mathsf{F}_{1}+\frac{2J}{D}\mathsf{F}_{2}$ (29)
@@
    then we can define $\mathsf{F}_{1,X+\bar{X}}\equiv\sum_{u}c_{u}\mathsf{F}_{0,X+\bar{X}}$. Notice
@@
We find that three new graph categories, $\mathsf{F}_{2},\mathsf{F}_{1,X+\bar{X}},\mathsf{F}_{0,X\bar{X}}$, emerge when calculating $\mathcal{L}_{2}^{2}$. Next, we consider the action of $\mathcal{L}_{2}$ on these new categories.
@@
  • $\mathcal{L}_{2}\mathsf{F}_{2}$: Notice that only the first term in Eq. (22) survives. The action of $\mathcal{L}_{2}$ on $\mathsf{F}_{2}$ is straightforward to obtain
@@
  • $\mathcal{L}_{2}\mathsf{F}_{1,X+\bar{X}}$: Using Eq.(22), we have
@@
  • $\mathcal{L}_{2}\mathsf{F}_{0,X\bar{X}}$: Using Eq.(22), we have
@@
Here, we find two new graph categories, $\mathsf{F}_{2,X+\bar{X}}$ and $\mathsf{F}_{1,X\bar{X}}$. We will then study the action of $\mathcal{L}_{2}$ on these categories.
@@
  • $\mathcal{L}_{2}\mathsf{F}_{2,X+\bar{X}}$: Using Eq.(22), we have
@@
  • $\mathcal{L}_{2}\mathsf{F}_{1,X\bar{X}}$: Using Eq.(22), we have
@@
    $=\left(w+J\right)\mathsf{F}_{1,X\bar{X}}+\frac{2J}{D}\mathsf{F}_{2}$
@@
Rank $r$ New Categories $\mathsf{F}$
@@ 4 $\emptyset$ 0
So far, we have found that no new categories have appeared. In summary, we list the total of 8 categories in Table 2.
@@ 3.2 Observables
$\begin{bmatrix}f_{1}(t)\\ f_{2}(t)\\ f_{3}(t)\\ f_{4}(t)\\ f_{5}(t)\\ f_{6}(t)\\ f_{7}(t)\\ f_{8}(t)\end{bmatrix}=\begin{bmatrix}0&0&\frac{1}{4}&\frac{1}{2}&\frac{1}{4}\\ 0&\frac{D^{2}-2}{D\left(D^{2}-4\right)}&\frac{-1}{4(D-2)}&\frac{-1}{2D}&\frac{-1}{4(D+2)}\\ 0&0&-\frac{1}{4}&0&\frac{1}{4}\\ 0&\frac{-1}{D^{2}-4}&\frac{1}{4(D-2)}&0&\frac{-1}{4(D+2)}\\ \frac{1}{D^{2}-1}&\frac{-2}{D^{2}-4}&\frac{1}{2(D-1)(D-2)}&0&\frac{1}{2(D+1)(D+2)}\\ 0&0&\frac{1}{4}&-\frac{1}{2}&\frac{1}{4}\\ \frac{-1}{D^{3}-D}&\frac{4}{D(D^{2}-4)}&\frac{-1}{2(D-1)(D-2)}&0&\frac{1}{2(D+1)(D+2)}\\ 0&\frac{2}{D(D^{2}-4)}&\frac{-1}{4(D-2)}&\frac{1}{2D}&\frac{-1}{4(D+2)}\end{bmatrix}\times\begin{bmatrix}1\\ e^{-Jt}\\ e^{-(2-2D)^{-1}Jt}\\ e^{-2Jt}\\ e^{-(2+2D)^{-1}Jt}\end{bmatrix}$ (52)
Generally, we can calculate
$\text{Tr}\left(O_{1}(t)O_{2}O_{3}(t)O_{4}\right)=\hbox{$\vbox{\hbox{\resizebox{}{}{{\hbox{\set@color{}}}}}}$}\equiv\sum_{a=1}^{8}f_{a}(t)\times\left(e^{-\text{i}\mathsf{E}t}\mathsf{F}_{a}\right)_{O_{1},O_{2},O_{3},O_{4}}\ .$ (53)
The noise-free case $J=0$, all $f_{a}(t)$ vanishes except $f_{1}(t)=1$, one can find
@@
which is just the unitary time evolution for a closed system. Using Eq.(53), we can calculate the fluctuations of two-point functions, spectral form factor (SFF) and other obervables.
There is an alternative method Guo:2024zmr to evaluate $\mathcal{U}_{2}(t)$ by observing that
@@
[…]